Slope two bodies with the same m = 1kg mass in relation to weightless wires
and put one on the turntable that was at the top of the plane,
and one on an inclined plane. The inclined plane is horizontal
Create an angle α = 30o. If the body on an inclined plane slides without friction
Find the acceleration and pulling force of each body. Turns back without rubbing
1
Expert's answer
2020-03-09T11:18:01-0400
"ma=T"

"ma=mg\sin{30}-T=0.5mg-T"

So,


"2ma=0.5mg"

"a=0.25g=0.25(9.8)=2.45\frac{m}{s^2}"

The pulling force of each body:

"F=ma=(1)(2.45)=2.45\ N"